OpenStudy (anonymous):

Rewrite the following equation as a function of X : 23x-1/4y=41=0

OpenStudy (anonymous):

move everything that does NOT have Y to the right side. then isolate Y to get a function of X.

OpenStudy (anonymous):

why are there 2 equal signs?

OpenStudy (anonymous):

oops it supposed to be a minus sign sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the 1/4y = (1/4y) right? or do you mena something else

OpenStudy (anonymous):

Where do we place the 23x yeah it's a fraction

OpenStudy (anonymous):

the 23x goes to the right side

OpenStudy (anonymous):

subtract 23x from 0?

OpenStudy (anonymous):

23x-(1/4y)-41=0. -(1/4y) = 41-23x. (1/4y)=-41+23x. 1 = (-41+23x)(4y). 4y= 1/(-41+23x) . y = [1/(-41+23x)]/4.

OpenStudy (anonymous):

sorry if it looks confusing. yes it is 0 -23x. so that becomes NEGATIVE 23x

OpenStudy (anonymous):

the steps go left to right not top to bottom

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

no problem! and for the 3rd line of the steps, i multiplied each side by 4y.